OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line through point (-3, 4) and perpendicular to a line that has slope 2/5 =?

OpenStudy (anonymous):

Is my answer 5x+2y=y=7?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

to get the equation of the line use formula y-y1 = m(x-x1) for m use -5/2

OpenStudy (anonymous):

m = -5/2 y1 = 4 x1 = -3 are all the figures u need can u plug them into the formula nd work it out?

OpenStudy (anonymous):

do u understand where the m = -5/2 came from?? or u want me to xplain??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol yes to which sorry wasnt specific enough :)